# Positive microlocal holonomies are globally regular

### v1

@@ A.2. Dg-categories and sheaves in contact and symplectic topology
In part due to historical reasons, the scientific development of contact and symplectic topology with regards to the study of Legendrian submanifolds has been more inclined towards the framework of dg-categories, and that of $A_{\infty}$-categories, a close relative of dg-categories. (In a sense, $A_{\infty}$-categories form minimal models for dg-categories.) To wit, the Legendrian contact dg-algebra [Che02, EES05] and various Fukaya categories, cf.  [Sei08, Chapter 2] and [GPS20, Section 3], naturally inherit their dg- and $A_{\infty}$-structures via the geometric counts extracted from certain moduli spaces of pseudo-holomorphic disks. The $\operatorname{Aug}_{+}$ and $\operatorname{Aug}_{-}$ $A_{\infty}$-categories, respectively studied in [NRS+20] and [BC14], are also instances of this phenomenon. If we study symplectic topology via the microlocal theory of $\infty$-sheaves, with coefficients in an $\infty$-category, a first dissonance appears at the algebraic level: Floer theoretic constructions produce dg-structures or $A_{\infty}$-structures, rather than $\infty$-structures. There are ways to address this:
@@
It might be possible to establish results for derived stacks as in [TV07] for certain compactly-generated presentable $k$-linear stable $\infty$-categories, instead of finite-type dg-categories. See e.g. [BZFN10, Lur18, GR17a, GR17b] for the formalism of derived stacks using stable $\infty$-categories and [Pan11], [Low12, Section 4] and [PT22, Section 7] for some construction and results of derived moduli stacks (partially) in the setting of $\infty$-categories. $\Box$
@@
[…]

### v2

@@ A.2. Dg-categories and sheaves in contact and symplectic topology
In part due to historical reasons, the scientific development of contact and symplectic topology with regards to the study of Legendrian submanifolds has been more inclined towards the framework of dg-categories, and that of $A_{\infty}$-categories, a close relative of dg-categories. (In a sense, $A_{\infty}$-categories form minimal models for dg-categories.) To wit, the Legendrian contact dg-algebra [CHE02, EES05] and various Fukaya categories, cf.  [SEI08, Chapter 2] and [GPS20, Section 3], naturally inherit their dg- and $A_{\infty}$-structures via the geometric counts extracted from certain moduli spaces of pseudo-holomorphic disks. The $\operatorname{Aug}_{+}$ and $\operatorname{Aug}_{-}$ $A_{\infty}$-categories, respectively studied in [NRS+20] and [BC14], are also instances of this phenomenon. See [COS19, COS24, PAS24, TAN24] for more details on the relation between dg-categories and $A_{\infty}$-categories. If we then study symplectic topology via the microlocal theory of $\infty$-sheaves, with coefficients in an $\infty$-category, a first dissonance appears at the algebraic level: Floer theoretic constructions produce dg-structures or $A_{\infty}$-structures, rather than $\infty$-structures. There are ways to address this:
@@
It might be possible to establish results for derived stacks as in [TV07] for certain compactly-generated presentable $k$-linear stable $\infty$-categories, instead of finite-type dg-categories. See e.g. [BFN10, LUR18, GR17a, GR17b] for the formalism of derived stacks using stable $\infty$-categories and [PAN11], [LOW12, Section 4] and [PT22, Section 7] for some construction and results of derived moduli stacks (partially) in the setting of $\infty$-categories. $\Box$
@@
This rectification is summarized as follows. Namely, [COH16, Cor. 5.7] shows that the underlying $\infty$-category $\mbox{dg-cat}_{k}$ of the Morita model category structure on the category of small dg-categories is equivalent to the $\infty$-category of compactly-generated presentable $k$-linear stable $\infty$-categories $\operatorname{Pr}^{\text{L}}_{\omega,st,k}$ (with functors that preserve colimits and compact objects) or equivalently $(\operatorname{Pr}^{\text{R}}_{\omega,st,k})^{op}$ (with functors that preserve colimits and limits). At core, the results of [COH16] are rectification results, in particular showing that given a compactly generated $k$-linear stable $\infty$-category, there exists a pretriangulated dg-category corresponding to it, in the appropriate homotopical sense. (See also the rectification result [HAU15, Theorem 1.1], which implies that $\infty$-categories enriched over chain complexes are equivalent to dg-categories.)